The weight of a rectangular block of metal varies jointly as its length, width and thickness. if a weight of a 12-in by 8-in by 6-in block of aluminum is 18.7 pounds, find the weight of a 16-in by 10-in by 4-in block.
I have ni clue how to solve it. I would appreciate any help

Answer by mananth(16949)

12-in by 8-in by 6-in block of aluminum is 18.7 pounds,
Volume = 12*8*6 = 576 in^3
weight = 18.7
density = mass /volume
density = 18.7/576
density = 0.032 lb/cu.in
find the weight of a
Volume of block 16-in by 10-in by 4-in block.
Volume = 640 cu.in
Mass = Density * volume
Mass = 0.032*640
Mass = 20.78 pounds
m.ananth@hotmail.ca
I have ni clue how to solve it.